# BAPO: Base-Anchored Preference Optimization for Personalized Alignment in Large Language Models

### v1

@@ Abstract
While learning to align Large Language Models (LLMs) with human preferences has shown remarkable success, aligning these models to meet the diverse user preferences presents further challenges in preserving previous knowledge. This paper examines the impact of personalized preference optimization on LLMs, revealing that the extent of knowledge loss varies significantly with preference heterogeneity. Although previous approaches have utilized the KL constraint between the reference model and the policy model, we observe that they fail to maintain general knowledge and alignment when facing personalized preferences. To this end, we introduce Base-Anchored Preference Optimization (BAPO), a simple yet effective approach that utilizes the initial responses of reference model to mitigate forgetting while accommodating personalized alignment. BAPO effectively adapts to diverse user preferences while minimally affecting global knowledge or general alignment. Our experiments demonstrate the efficacy of BAPO in various setups.
BAPO: Base-Anchored Preference Optimization for Personalized
Alignment in Large Language Models
@@ 1 Introduction
Figure 1: Overview of Base-Anchored Preference Optimization (BAPO): For a given user prompt, the base response achieves general alignment. Models A and Model C, fine-tuned with BAPO, maintain this alignment by anchoring to the base response. In contrast, Model B, fine-tuned with DPO, fails to preserve the knowledge from the base response, drifting away from the desired knowledge preservation area. The full example is provided in Appendix E
@@ 3.2 Theoretical Analysis
The proof is provided in Appendix D. This proposition suggests that the last $k$-dimensional subspace, which governs personalized rewards, is typically much smaller than the first $d-k$-dimensional subspace responsible for general alignment. Consequently, this reduction in complexity of the parameter space influencing personalized rewards allows for more efficient estimation with fewer samples. In practice, personalization is often driven by a smaller, critical set of features compared to those affecting broader alignment criteria.
@@
The results in Figure 6 empirically support our claim. We analyze the changes in log probabilities of responses under DPO and BAPO. For experimental details, see section 4. While log probabilities for the base response decrease under DPO, BAPO maintains them, enhancing stability throughout the preference optimization process. This consistency enables the model to assign significantly higher log probabilities to the chosen response, thus speeding up the learning process.
@@ 5 Related Work
The issue of forgetting previously acquired knowledge during the fine-tuning on heterogeneous data has been extensively discussed (Wang et al., 2023b, 2024b), highlighting a fundamental trade-off between preserving old knowledge and acquiring new knowledge (Parisi et al., 2019; McCloskey and Cohen, 1989). In the context of LLMs, recent research has shown that the knowledge from pre-training can be compromised by supervised fine-tuning (SFT) on instruction data (Dou et al., 2023; Dong et al., 2023). A similar issue, known as alignment tax, occurs in preference optimization (Lin et al., 2023; Lu et al., 2024). While prior research has addressed the forgetting induced by the LLM fine-tuning (Biderman et al., 2024; Luo et al., 2023), the effects of accommodating specific user preferences and the impact of their heterogeneity remain largely unexplored. In our study, we investigate how personalized preference optimization affects both global knowledge and general alignment.

### v2

# BAPO: Base-Anchored Preference Optimization for Overcoming Forgetting in Large Language Models Personalization

@@ Abstract
While learning to align Large Language Models (LLMs) with human preferences has shown remarkable success, aligning these models to meet the diverse user preferences presents further challenges in preserving previous knowledge. This paper examines the impact of personalized preference optimization on LLMs, revealing that the extent of knowledge loss varies significantly with preference heterogeneity. Although previous approaches have utilized the KL constraint between the reference model and the policy model, we observe that they fail to maintain global knowledge and general alignment when facing personalized preferences. To this end, we introduce Base-Anchored Preference Optimization (BAPO), a simple yet effective approach that utilizes the initial responses of reference model to mitigate forgetting while accommodating personalized alignment. BAPO effectively adapts to diverse user preferences while minimally affecting global knowledge or general alignment.
BAPO: Base-Anchored Preference Optimization for Overcoming Forgetting in Large Language Models Personalization
@@ 1 Introduction
Figure 1: Overview of Base-Anchored Preference Optimization (BAPO): For a given user prompt, the base response achieves general alignment. Models A and Model C, fine-tuned with BAPO, maintain this alignment by anchoring to the base response. In contrast, Model B, fine-tuned with DPO, fails to preserve the knowledge from the base response, drifting away from the desired knowledge preservation area. The full example is provided in Appendix E.
@@ 3.2 Theoretical Analysis
The proof is provided in Appendix D. This proposition implies that the $k$-dimensional subspace, which governs personalized rewards, is typically much smaller than the $(d-k)$-dimensional subspace responsible for general alignment. Consequently, this reduction in the complexity of the parameter space influencing personalized rewards allows for more efficient estimation with fewer samples. In practice, personalization is often driven by a smaller, critical set of features compared to those affecting broader alignment criteria.
More intuitively, the theoretical analysis in Proposition1 suggests that identifying the parameter dimensions crucial for global knowledge and general alignment (${\theta}_{G}$) can enhance the process of optimizing personalized preferences. This is achieved by focusing on the parameter dimensions essential for personalization (${\theta}_{L}$), under the assumption that ${\theta}_{G}$ and ${\theta}_{L}$ are separable. The results in Figure 6 empirically support our claim. We analyze the changes in log probabilities of responses under DPO and BAPO. The experimental details are provided in section 4. While the log probabilities for the base response decrease under DPO, BAPO maintains them, enhancing stability throughout the preference optimization process. This consistency enables the model to assign significantly higher log probabilities to the chosen response, thus accelerating the learning process.
@@ 4.4 Ablation Study
5 Discussion
We demonstrate that BAPO benefits from the separation between general alignment and personalized preferences. However, preserving the likelihood of the base response may sometimes hinder alignment with personalized preferences. For example, if a user prefers humorous but potentially harmful responses, and the base response is harmless, maintaining the base likelihood can conflict with achieving the user’s preferred style. Thus, while we distinguish between ’Global Knowledge & General Alignment’ and ’Personalized Alignment,’ their interaction is influenced by both the model’s existing knowledge and the new, targeted personalization.
Our experiments show that this separation is advantageous for BAPO, as personalized preferences—whether in specific response styles (P Soups Datasets) or specific domains (DSP Datasets)—tend not to overlap significantly with the model’s general knowledge. Nevertheless, the challenge arises when preserving the base response’s likelihood ($y_{b}$) compromises the ability to cater to personalized preferences. This situation parallels the Stability-Plasticity Dilemma in conventional Continual Learning (Mermillod et al., 2013; Wang et al., 2023b), where the relationship between old and new knowledge influences the potential for forgetting.
@@ 6 Related Work
The issue of forgetting previously acquired knowledge during the fine-tuning on heterogeneous data has been extensively discussed (Wang et al., 2023c, 2024b), highlighting a fundamental trade-off between preserving old knowledge and acquiring new knowledge (Parisi et al., 2019; McCloskey and Cohen, 1989). In the context of LLMs, recent research has shown that the knowledge from pre-training can be compromised by supervised fine-tuning (SFT) on instruction data (Dou et al., 2023; Dong et al., 2023). A similar issue, known as alignment tax, occurs in preference optimization (Lin et al., 2023; Lu et al., 2024). While prior research has addressed the forgetting induced by the LLM fine-tuning (Biderman et al., 2024; Luo et al., 2023), the effects of accommodating specific user preferences and the impact of their heterogeneity remain largely unexplored. In our study, we investigate how personalized preference optimization affects both global knowledge and general alignment.
@@ 7 Conclusion
Acknowledgements
This work was supported by Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government(MSIT) [No. RS-2024-00457882, AI Research Hub Project, 90%] and [No. 2019-0-00075, Artificial Intelligence Graduate School Program (KAIST), 10%].
@@ Acknowledgements
Use of AI Assistants
We utilize Copilot [^1] https://github.com/features/copilot for the development of our code pipeline, primarily for its auto-completion capabilities. For drafting the paper, we employ ChatGPT [^2] https://chatgpt.com/ to review the content, focusing on identifying grammatical errors and awkward expressions. However, the core content of the paper is original. We do not rely on AI assistants to generate any specific content that is closely related to the main claims of our research.